You run a school in Florida. Fixed monthly cost is $5,953.00 for rent and utilities, $6,358.00 is spent in salaries and $1,076.00 in insurance. Also every student adds up to $103.00 per month in You are considering moving the school to another neighborhood where the rent and utilities wil increase to $11,119.00, salaries to $6,518.00 and insurance to $2,174.00 per month. Variable cost per student will increase up to $182.00 per month. However you can charge $1,073.00 per student. At what point will you be indifferent between your current mode of operation and the new option?
Answer #1

Answer:

let x be the number of students

So

Profit for existing location=x*Per month student charge-x*per month stationary etc-Rent & utility-Salaries-insurance

Profit for existing location=x*711-x*103-5953-6358-1076=608x-13387

Profit for New location=x*Per month student charge-x*per month stationary etc-Rent & utility-Salaries-insurance

Profit for New location=x*1073-x*182-11119-6518-2174=891x-19811

For indifferent between new location and existing location

Profit for existing location=Profit for New location

608x-13387=891x-19811

x=22.69(23 students)